OpenStudy (superhelp101):

What are the possible rational zeros of f(x) = x4 - 4x3 + 9x2 + 5x + 14? ± 1, ± 1 over 2, ± 1 over 7, ± 1 over 14 ± 1, ± 2, ± 7, ± 14 ± 1, ± 1 over 4,± 1 over 5, ± 1 over 9, ± 1 over 14 ± 1, ± 4, ± 5, ± 9, ± 14

OpenStudy (superhelp101):

can someone explain how to do this? :/

OpenStudy (superhelp101):

@vishweshshrimali5 @tkhunny @ganeshie8 @Compassionate @dumbcow :)

OpenStudy (anonymous):

well theres going to 4 answers

OpenStudy (superhelp101):

yes

OpenStudy (anonymous):

so c and d is out

OpenStudy (anonymous):

dumcow will tell you the rest

OpenStudy (dumbcow):

its the factors of last coefficient over factors of first coefficient

OpenStudy (tkhunny):

Please stop tagging the world. You will have to read the "Rational Root Theorem". Is this familiar?

OpenStudy (vishweshshrimali5):

Is your function correct ?

OpenStudy (anonymous):

1 is not a root

OpenStudy (vishweshshrimali5):

It appears to be incorrect

OpenStudy (superhelp101):

it doesn't make much sense. I have read it. I understand how that we need 4 answers.

OpenStudy (vishweshshrimali5):

Due to what @RaphaelFilgueiras said, I suspect that you did some mistake in writing the function

OpenStudy (superhelp101):

that is exacting what my question states :/

OpenStudy (vishweshshrimali5):

This function has no real root

OpenStudy (superhelp101):

that's a little odd...

OpenStudy (vishweshshrimali5):

Well it appears that there *is* some mistake.

OpenStudy (anonymous):

hmmmm

OpenStudy (tkhunny):

It asks for possible rational roots. This is understood to be a priori and the existence of actual rational roots is of no consequence. The correct answer is given in the choices. There is no observable error.

OpenStudy (vishweshshrimali5):

But there is no possible rational root @tkhunny

OpenStudy (anonymous):

so its between answers A and B

OpenStudy (vishweshshrimali5):

What does possible rational root mean ?

OpenStudy (vishweshshrimali5):

Because using \(\pm\) would mean 2 solutions. Each option would give 8 solutions

OpenStudy (dumbcow):

ahh this theorem is becoming obsolete now with graphing calculators and wolfram

OpenStudy (superhelp101):

this was in my lesson, I don't get this: To find all of the possible rational zeros of the polynomial function, divide each factor of p by each factor of q. In other words, find all possible combinations of p over q. p over q = ±1 over 1, ±2 over 1, ±3 over 1, ±6 over 1 When each of the fractions is simplified, the possible rational zeros are found: ±1, ±2, ±3, ±6 Keep in mind that ±1, ±2, ±3, and ±6 are only the possible rational zeros. Irrational and complex zeros cannot be identified by the p over q values. Also, not all of these eight possible zeros will actually be zeros. Remember, the Fundamental Theorem of Algebra tells us that the function has exactly five zeros. To find the zeros of the function, test each of the possible zeros in the synthetic division process to discover which results in a remainder of 0.

OpenStudy (vishweshshrimali5):

Ohh now I get it.. Sorry about my all above comments :)

OpenStudy (vishweshshrimali5):

But what are p and q ?

OpenStudy (superhelp101):

It's ok... from the lesson this is what they are referring to....f(x) = 1x5 + 2x4 – 4x3 – 4x2 – 5x – 6

OpenStudy (tkhunny):

1) Look at an equation. 2) Determine the possible numbers of Positive and Negative Real roots. 3) Determine the possible Rational Roots. 4) Combine the information from 2) and 3) to narrow down the list a little. 5) Actually search for rational roots. 6) Actually search for Real Roots. 7) Finish up by finding Complex Roots. What the problem statement is talking about is Step #3. In Step #3, we do not yet know that there are no Rational Roots. We are just listing possibilities - where we should look if there happen to be some. Later, we may discover that none actually IS a root, but we don't care about that in Step #3 - we don't know enough, yet.

OpenStudy (tkhunny):

From the lesson, "Also, not all of these eight possible zeros will actually be zeros. "

OpenStudy (superhelp101):

that is what I am stuck on...

OpenStudy (vishweshshrimali5):

Well we can find out the step 2 answer using Descarte's theorem. But I think as @tkhunny said, the question is asking about step 3

OpenStudy (vishweshshrimali5):

In the lesson it says "divide each factor of p by each factor of q. In other words, find all possible combinations of p over q." Are p and q some coefficients or some functions or anything else ?

OpenStudy (vishweshshrimali5):

Were you given anything related to p and q in the lesson ?

OpenStudy (anonymous):

i belive that you need to do are look at the factors of the constant 14...q ±1,±2,±7,±14 and the coefficient of the leading term which is ±1 and the possible rational roots can be qp

OpenStudy (superhelp101):

It said the factors of P=+-1, +-2, +-3, +-6 and the factors of q: +-1

OpenStudy (superhelp101):

alright.. I guess I'll try to figure this one out on my own. thank you for your suggestions guys. :)

OpenStudy (anonymous):

i believe its B though

OpenStudy (superhelp101):

;)

OpenStudy (dumbcow):

factors of 14: 1,2,7,14 factors of 1: 1 --> +- 1/1 ,2/1 , 7/1, 14/1 its not hard once you grasp the concept

OpenStudy (tkhunny):

What are you still figuring out? f(x) = 1x5 + 2x4 – 4x3 – 4x2 – 5x – 6 In standard form, The leading coefficient is 1. All the factors of 1 are 1. That's easy. These are your q's. The constant term is -6. All the factors of 6 are 1, 2, 3, 6. That's easy. These are your p's. Now, we just make a collection of all combinations. Since q presents only 1, we will get only integers. +/-1, +/-2, +/-3, +/-6 -- Done

OpenStudy (tkhunny):

Slightly more complicated: (x) = 2x5 + 2x4 – 4x3 – 4x2 – 5x – 8 In standard form, The leading coefficient is 2. All the factors of 2 are 1, 2. That's easy. These are your q's. The constant term is -8. All the factors of 8 are 1, 2, 4, 8. That's easy. These are your p's. Now, we just make a collection of all combinations. +/-1, +/-2, +/-4, +/-8 -- Using q = 1 +/-1/2, +/-2/2, +/-4/2, +/-8/2 -- Using q = 2 There is some redundancy, since 2/2 = 1 and we have that already. 4/2 = 2 and we have that already. 8/2 = 4 and we have that already. The only new suspect we found with q = 2 is +/- 1/2.

OpenStudy (superhelp101):

how about for this equation f(x) = x4 - 4x3 + 9x2 + 5x + 14?

OpenStudy (tkhunny):

You tell me. Your numerators are factors of 14. Name them. Your denominators are factors of 1. Name them.

OpenStudy (dumbcow):

@superhelp101 , are you not seeing my posts??

OpenStudy (anonymous):

lol

OpenStudy (superhelp101):

yes I see until this part....but i am still not understanding the concepts of dividing p by q

OpenStudy (superhelp101):

relating to the problem i posted

OpenStudy (tkhunny):

Start simple. What is the root of this equation? 3x + 5 = 0?

OpenStudy (superhelp101):

I don't really get how and why you divided these +- 1/1 ,2/1 , 7/1, 14/1

OpenStudy (dumbcow):

forget p and q think last/first if that makes it easier

OpenStudy (tkhunny):

Solve my little linear problem. You will see it.

OpenStudy (anonymous):

rational root theorem

OpenStudy (superhelp101):

x=-5/3

OpenStudy (tkhunny):

How did that 5 get into the numerator? How did that 3 get into the denominator? That's ALL we're doing. If a root is going to be rational, it has to be the result of some linear factor that looks just like that. The only factors that can look like that are (Some Factor of the leading coefficient * x + Some factor of the Constant Term) 2x^2 +5x - 3 = (2x-1)(x+3) = 0 ==> x = 1/2 or x = -3/1 Notice how the numerators, 1 and 3, are factors of the 3 we had in the first place. Notice how the denominators, 2 and 1, are factors of the 2 we had in the first place. This is all we are doing. Playing with linear factors that you naturally do, anyway.

OpenStudy (superhelp101):

ok

OpenStudy (tkhunny):

You better show us another one. Is this problem set done? I am not yet filled with warm, fuzzy feelings.

OpenStudy (superhelp101):

it makes sense, if i have any questions i'll let you know ;)

OpenStudy (tkhunny):

Past my bedtime. I'll have to try to sleep while I'm worried about you. :-)

OpenStudy (superhelp101):

plz don't let me bother you, it's 1:30 am here.. past mine too

OpenStudy (anonymous):

what??!!!

OpenStudy (anonymous):

its 2:35 PM here

OpenStudy (superhelp101):

wow! I am a hard worker ;)

OpenStudy (anonymous):

hmmmm not really. : D

OpenStudy (anonymous):

jk

OpenStudy (superhelp101):

lol, are you near tokyo?

OpenStudy (anonymous):

pretty much but not in Japan nor China

OpenStudy (superhelp101):

where then?

OpenStudy (anonymous):

Korea

OpenStudy (superhelp101):

lol my best friend is from korea ;)